40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 947     Entry time: Wed Aug 18 18:15:49 2010
Author: Alastair, Zach and Rana 
Type: Laser 
Category: GYRO 
Subject: PDH locking of the CCW loop 

Looking at the error signal for locking the laser to the gyro (CCW loop) while we sweep the laser frequency we have the following plots.  Blue is error signal, yellow the reflection signal from the PD, pink is the drive signal and blue is the transmission signal.

The first shot show how the system was to begin.  Rana pointed out that the signal is messy on either side of resonance suggesting something not-so-nice is happening.  We diagnosed that the sidebands were partially leaking into the cavity on resonance.  The sideband frequency was 14MHz at this point.

We tried increasing the frequency to 20MHz but the signal got a lot smaller (due to bandwidth of the PDs).  Instead we turned it down to 6MHz and now get  the second plot where the error signal is nice and smooth and the sideband signals on either side are asymmetric.

 RA:   We also found that attaching a long-ish BNC cable to the PZT drive input (also used for test injections) causes the loop to oscillate at ~235 kHz. It must be that the extra capacitive load is the problem, but I'm not sure why there is any effect. Could be that its the PZT resonance, but that's unknown until Zach/Dmass do the laser beating PZT sweep measurements.

Attachment 1: TEK00003.PNG  14 kB  | Hide | Hide all
TEK00003.PNG
Attachment 2: TEK00004.PNG  14 kB  | Hide | Hide all
TEK00004.PNG
ELOG V3.1.3-